Summary of what happened during Optics data taken on April 6.
Run 2642
*LUMIS*
Lumi signals were all saturated. They were set to produce ~150K channels with 50uA on Thick Carbon with Septum OFF (HALOG entry 316140). These values were: 900,990,1900,1800,1280,1500,1130,920
From Rupesh's Lumi studies, summarized in HALOG entry 316729, when Septum was turned ON, the signals for most of the lumis dropped by factors of 1.2-3.9 (38-129K channels) while the signals for lumis 3 and 7 were saturated.
From this, we can see that the lumi signals were much higher than we would expect.
*DETECTORS*
Detector Voltages were set to 50uA on Thick Carbon voltages. These values were confirmed by looking at the 'End of Run' HALOG entry 316561:
L-HRS ... card 2 chan 2 = 720 (for det 4 = lower quartz) chan 4 = 750 (for det 3 = upper quartz)
R-HRS ... card 14 chan 4 = 750 (for det 2 = upper quartz) chan 6 = 850 (for det 1 = lower quartz)
The signals on thick carbon should be ~50K channels, according to HAPLOG entry 2106, but they are now ~10-12K channels. The optics target (at 210 mg combined) is about half the thickness of the thick carbon so we have less than half of the expected rate for this thickness of carbon. This could be because not all of the foils are in the detector acceptance and can be checked with our counting mode optics runs.
Run 2643
*LUMIS*
For this run, lumi voltages were set to random lower values to avoid saturation. By my best recollection, the voltages were: 700,700,1000,1000,700,1000,700,700
*DETECTORS*
Same as previous run.
Run 2644
*LUMIS*
Lumi voltages were set to values used for thick lead running before the run started. These values are: 600,590,760,850,740,755,555,530
From run 2664 (50uA on Thick Carbon), taken later that day, the Lumi signals for these voltages are ~5K channels (see Figure 1). In this run the signals start from ~40-120K channels and rise between 60-100% (see Figure 2), so for our optics target, half as thick, this corresponds to between 16 and 70 times the expected rate.
Figure 1
Figure 2